Mark is 10 years younger then Larry.Larry's age in 8 years will exceed twice Mark's age 3 years ago by 4 years.How old is each now?
Answer by stanbon(75887)

Mark is 10 years younger then Larry.
Larry's age in 8 years will exceed
twice Mark's age 3 years ago by 4 years.
How old is each now?
--------------------------
NOW DATA:
Let Larry's age be "x": Mark's age is "x-10"
----------
In 8 years years Larry will be x+8
x-13 (Mark;s age three years ago)
--------------
EQUATION:
x+8 = 2(x-13)+4
x+8 = 2x-22
x = 30 (Larry's age now)
x-10 = 20 (Mark's age now)
================
Cheers,
Stan H.